OpenStudy (anonymous):

Which is an example of the commutative property? A.–13 + 13 = 0 B.–(11 + 4) = –11 + (–4) C.18 + (–14) = –14 + 18 D.9 + (–4 + 7) = (9 + –4) + 7

OpenStudy (anonymous):

B -(11+4)= -11-4= -15 -11=(-4)=-11-4=-15